# Bézier Splatting for Fast and Differentiable Vector Graphics

### v2

@@ 4.2 Method Comparison
Table 1 reports the forward and backward runtime for processing 2,048 curves on an image with a resolution of 2,040×1,344. Compared to DiffVG [13], our method significantly accelerates VG rasterization by approximately 20$\times$ faster per forward step and 150$\times$ faster per backward step for open curves. For closed curves, our color filling strategy requires sampling 20 additional Bézier curves per closed curve, while our method remains highly efficient, achieving 6$\times$ faster forward and 28.5$\times$ faster backward computation.
@@
Table 2 quantitatively evaluates the quality of differentiable VG representations by three commonly used metrics, MS-SSIM [24], PSNR, and LPIPS [31]. Our method demonstrates superior performance in both training efficiency and image fidelity, under all curve numbers and open and closed curves. Our method achieves 6$\times$-13$\times$ training acceleration compared to DiffVG [13] and 68$\times$-90$\times$ compared to LIVE [16]
@@
Fig. 4 visually compares DiffVG [13], LIVE [3], and our Bézier splatting under varying numbers of open/closed curves. As the number of paths increases, it shows progressive enhancement in the visual details of mountains, lakes, and clouds. Compared to DiffVG [13], our method captures significantly more fine-grained textures. Compared to LIVE [16], our method effectively reduces artifacts in the areas of clouds and water, resulting in higher-fidelity rendering results. This improvement stems from our method’s ability to globally optimize all curves, rather than LIVE’s layer-by-layer path-adding strategy. By optimizing the entire VG simultaneously, our approach ensures a cohesive and natural appearance, avoiding the accumulation of errors and inconsistencies introduced by layer-wise updates.
@@
Table 3: Ablation study on adaptive pruning and densification strategy and the number of opacity parameters per curve (tested with 512 open curves on DIV2K dataset [22]).
@@
Full (w/ Prune&Densify, 3-opacity) 0.65 23.79 0.50
- w/o Prune&Densify 0.63 23.23 0.50
@@
As shown in Fig. 3, Bézier splatting consistently renders high-quality VGs across various image domains, from photorealistic natural images, watercolor paintings, to clip arts and cartoon images. Unlike learning-based image vectorization methods [14, 19, 1], which can not easily generalize to out-of-domain data, our method can flexibly handle different types of images.
4.3 Ablation study
Adaptive pruning and densification. We evaluate the effectiveness of our pruning and densification strategy and the quantitative results are shown in the third row of Table 3. The pruning step effectively removes redundant curves, while the densification step utilizes an error map derived from the loss function to initialize new curves, allowing for allocating more curves to areas with fine details. Fig. 5 also demonstrates the texture quality improvement brought by this strategy.
Multi-opacity strategy. To enhance the VG representation quality without significantly increasing the number of parameters, we assign opacity values to Bézier curves. For open curves, we assign a single opacity value to each Bézier curve. For closed areas, we adopt three opacity values and compute their linear interpolation across the enclosed region, assigning these values to the newly generated Bézier curves. As shown in Table 3, incorporating richer opacity variation improves the rendering fidelity. Without the multi-opacity strategy, as shown in the third image of Fig. 5, artifacts such as visible lines crossing objects appear (e.g., a line across the chair and doll). The multi-opacity strategy helps mitigate this issue by using opacity variations as an effective criterion for adaptively pruning the curves which have low opacities in between the enclosed regions, leading to cleaner and more visually coherent results.
5 Conclusion
[…]
Limitation and future work. The closed curves in Bézier splatting are enforced to have convex shapes using Xing loss [16] to prevent false interpolation results, which may slightly reduce the capacity of VG representations. Additionally, the closed curves require sampling more Gaussian points to represent area boundaries precisely, leading to slower forward-backward compared to open curves. An interesting future direction is to leverage the efficiency and differentiability of our approach for high-quality VG synthesis applications such as text-to-VG generation or text-based VG editing.
Appendix A More Ablation Study Results
We progressively increase the number of curves for vectorizing a watercolor image with numerous small spots. As the number of curves grows, our approach first reconstructs the foreground object, the bird, then gradually refines the smaller spots as shown in Fig 6. This demonstrates that our method prioritizes the primary structure before optimizing finer details, effectively distributing curves to balance global structure and local texture representation, thanks to our adaptive pruning and densification strategy.
Figure 6: A comparison of curve number ($N$) by Bézier splatting.
Appendix B More Comparisons
We evaluate the differentiable VG methods on more natural images. We quantitatively evaluate our method on another natural image dataset, Kodak [12], and the results are shown in Table 4. Fig. 7 shows more qualitative comparisons. In the second image, our method successfully captures finer details in the blanket, preserving intricate textures and subtle variations that are less accurately represented by other methods. This demonstrates the superior representation capacity of our Bézier Splatting, particularly for complex textures and high-frequency details while maintaining structural integrity.
@@ Appendix B More Comparisons
Table 4: Quantitative evaluation on the Kodak dataset [12] (256 to 2,048 open curves).
Method 256 512 1024 2048
$SSIM^{\uparrow}$ $PSNR^{\uparrow}$ $LPIPS^{\downarrow}$ $SSIM^{\uparrow}$ $PSNR^{\uparrow}$ $LPIPS^{\downarrow}$ $SSIM^{\uparrow}$ $PSNR^{\uparrow}$ $LPIPS^{\downarrow}$ $SSIM^{\uparrow}$ $PSNR^{\uparrow}$ $LPIPS^{\downarrow}$
DiffVG 0.601 23.43 0.535 0.645 24.70 0.495 0.699 26.04 0.439 0.762 27.56 0.360
Ours 0.679 26.18 0.457 0.743 27.90 0.383 0.797 29.24 0.310 0.793 28.75 0.300
Appendix C More Rendering Results on Different Image Domains
@@
We evaluate our method on more image examples across a broader range of image domains. Fig. 8 shows more natural image examples by open curves, while Fig. 9 shows more results by closed curves. Fig. 10 shows more results on animation images. The results demonstrate that both the global structure and local texture of the images are well reconstructed, highlighting the effectiveness of our approach in capturing fine details and complex shapes.
@@
Figure 8: More rendering results by Bézier splatting on DIV2K dataset [22].
Figure 9: More rendering results by Bézier splatting on DIV2K dataset [22].
Figure 10: More rendering results by Bézier splatting on DanbooRegion dataset [29].

### v3

@@ 4.2 Method Comparison
Table 1 reports the forward and backward runtime for processing 2,048 curves on an image with a resolution of 2,040$\times$1,344. Compared to DiffVG [14], our method significantly accelerates VG rasterization by 31.4$\times$ faster per forward step and 149.2$\times$ faster per backward step for open curves. For closed curves, our color filling strategy requires sampling 20 additional Bézier curves per closed curve, while our method remains highly efficient, achieving 6$\times$ faster forward and 18.2$\times$ faster backward computation.
@@
Table 2 quantitatively evaluates the quality of differentiable VG representations by three commonly used metrics, MS-SSIM [25], PSNR, and LPIPS [33]. Our method demonstrates higher optimization efficiency and rendering fidelity for both open and closed curves with different curve numbers.
@@
Fig. 3 visually compares DiffVG [14], LIVE [3], [26], and our Bézier Splatting under 512 closed curves and 2048 open curves. Compared to DiffVG [14], our method captures significantly more fine-grained textures. Compared to LIVE [17], our method effectively enhances rendering quality, resulting in higher-fidelity results. This improvement stems from our method’s ability to globally optimize all curves, rather than LIVE’s layer-by-layer path-adding strategy. Compared to LIVSS [26], our method demonstrates superior ability in preserving fine structural details, especially in text regions. Due to its heavy reliance on layer-wise semantic simplification, LIVSS struggles to optimize regions where semantic information is ambiguous or difficult to extract. Moreover, by optimizing the entire VG simultaneously, our approach ensures a cohesive and natural appearance, avoiding the accumulation of errors and inconsistencies introduced by layer-wise updates.
As shown in Fig. 4, Bézier Splatting consistently renders high-quality VGs across various image domains from photorealistic natural images, watercolor paintings, to cartoon images. Unlike learning-based image vectorization methods [15, 20, 1], which can not easily generalize to out-of-domain data, our method can flexibly handle different types of images.
Figure 5: As a differentiable VGs renderer, Bézier Splatting can integrate with topology-aware strategies such as layer-wise vectorization [17] to further improve compositionality and detail.
Figure 6: Our Bézier Splatting representation is compatible with the standard SVG XML format and supports flexible vector editing operations.
4.3 Adaptability with Layer-wised Image Vectorization
Existing works such as LIVE [17], SGLIVE [36] and LIVSS [26] have demonstrated that layer-wise vectorization strategies can significantly improve the topology and compositionality of DiffVG-generated vector graphics [14]. Since the proposed Bézier Splatting is a fast and differentiable VGs renderer alternative to DiffVG, is naturally compatible with these topology-aware strategies and can leverage them to further enhance the structural quality of the optimized vector outputs. As shown in Fig. 5, we incorporate the layer-wise vectorization approach from LIVE [17] into the Bézier Splatting optimization process This integration yields improved compositional integrity and finer details for regions with complex patterns such as the butterfly’s wings.
4.4 Editability and Compatibility with SVG XML Format
Our Bézier Splatting representation is fully compatible with the standard SVG XML format, ensuring the same editability as standard SVGs and seamless integration with existing SVG tools. As shown in Fig. 6, a high-fidelity SVG can be obtained via a simple conversion algorithm (see Appendix for details). Once converted to SVG, the vectorized output is fully editable at the primitive level, allowing users to freely deform, remove, move, replicate, and recolor individual elements.
5 Conclusion
This work has presented Bézier Splatting, a novel differentiable vector graphics (VGs) representation that leverages Gaussian splatting for efficient Bézier curve optimization. Our method achieves 30$\times$ faster forward computation and 150$\times$ faster backward computation in rasterization compared to existing methods, while also delivering high rendering fidelity. Additionally, our adaptive pruning and densification strategy improves optimization by dynamically adjusting curve placement during optimization. Extensive experiments have demonstrated that Bézier Splatting outperforms existing differentiable VG methods in both training efficiency and visual quality, making it a promising solution for scalable applications of VGs.
Limitation and future work. The closed curves in Bézier Splatting are enforced to have convex shapes using Xing loss [17] to prevent false interpolation results, which may slightly reduce the capacity of VG representations. Additionally, the closed curves require sampling more Gaussian points to represent area boundaries precisely, leading to slower computation compared to open curves. An interesting future direction is to leverage the efficiency and differentiability of our approach for high-quality VG synthesis applications such as text-to-VG generation or text-based VG editing.
Appendix A Ablation Study
The effectiveness of adaptive pruning and densification and layer-wise vectorization strategy. We conduct an ablation study to evaluate the effectiveness of our adaptive pruning and densification strategy. Quantitative results are shown in Table 3. To further investigate the flexibility of our framework, we incorporate a layer-wise curve addition strategy inspired by LIVE [17], where curves are progressively added during training. Although this strategy achieves slightly better reconstruction metrics, it doubles the training time compared to our adaptive pruning and densification approach, making it less favorable for time-sensitive applications. Fig. 7 shows our result with layer-wise image vectorization strategy.
@@ Appendix A Ablation Study
Table 3: An ablation study on adaptive pruning and densification and layer-wise training strategy, evaluated on 512 closed curves with DIV2K dataset [23].
Method $SSIM^{\uparrow}$ $PSNR^{\uparrow}$ $LPIPS^{\downarrow}$ $Opt^{\downarrow}$
No Strategy 0.590 21.10 0.530 8.3 min
– w/ Prune&Densify 0.607 22.11 0.528 8.3 min
– w/ Layer-wise 0.613 22.21 0.521 16.2 min
Figure 7: Our Bézier Splatting is fully compatible with the layer-wise vectorization strategies [17].
Comparing different numbers of curves. We progressively increase the number of curves for vectorizing a watercolor image (Fig. 8) with numerous small spots. As the number of curves grows, our approach first reconstructs the foreground object, the bird, then gradually refines the smaller spots. This demonstrates that our method prioritizes the primary structure before optimizing finer details, effectively distributing curves to balance global structure and local texture representation, thanks to our adaptive pruning and densification strategy.
Figure 8: A comparison of different curve numbers by Bézier Splatting.
A systematic evaluation of computation time. To systematically evaluate the computational efficiency of our framework, we report detailed timing statistics across varying configurations of interpolation curve numbers and curve resolutions, as summarized in Table 4. The total forward time includes two major parts: the Gaussian sampling and the Gaussian splatting. Increasing the number of interpolated curves within the closed Bézier curves leads to a near-linear growth in both forward and backward runtimes, demonstrating good computational scalability. When fixing the interpolation number (e.g., at 40) and varying the number of Bézier curves, the forward time increases moderately, reflecting the additional cost of handling finer spatial detail. The splatting stage dominates the forward time at higher number of curves due to the increased number of sampled points per region, while the Gaussian sampling time is the primary bottleneck at lower number of curves. This is attributed to overheads from sequential memory allocation and data transfer operations, which are not effectively parallelized on the GPU, suggesting potential for further optimization.
Table 4: A systematic evaluation of computation time. We test the forward and backward time (ms) under varying numbers of sampled curves within one closed Bézier curve (Inter. #), and the total number of Bézier curves (Curve #).
Inter. # Curve # Forward w/ gradient Sample Gaussians Splatting Backward Render (FPS)
20 2048 10.66 4.12 6.11 21.64 103.45
40 2048 14.79 5.11 9.32 25.06 68.30
80 2048 27.01 8.22 17.74 33.40 37.90
40 256 7.20 3.81 3.09 16.72 150.30
40 512 7.70 3.88 3.67 17.42 144.30
40 1024 10.39 4.06 5.96 19.96 114.30
40 2048 14.79 5.11 9.32 25.06 68.30
Appendix B Convert Bézier Splatting to standard SVG Format
As described in Algorithm 1, we convert our optimized Bézier Splatting representation into a standard SVG file for compatibility with downstream vector editing tools. Each curve is represented by a set of $13k+$ 1 control points $\mathcal{C}_{i}\in\mathbb{R}^{(3k+1)\times 2}$, corresponding to $k$ continuous cubic Bézier segments. These control points are normalized to the range $[-1,1]$ and are first transformed into pixel coordinates according to the canvas size $(W,H)$.
For each curve $i$, we initialize the SVG path string $d$ with a [M $P_{i}^{0}$] command. We then iteratively construct each cubic segment using three internal control points: $(p_{1},p_{2},p_{3})=(P_{i}^{3j+1},P_{i}^{3j+2},P_{i}^{3j+3})$, and append the segment as a path command [C $p_{1}$, $p_{2}$, $p_{3}$] to $d$. After all segments are processed, a [Z] is appended to close the path.
Each path $d$ is then associated with a fill color $(r,g,b)$, computed via sigmoid activation on the optimized feature vector $\mathcal{F}_{i}$, and opacity 1.0. The result is stored in a global path set $\mathcal{S}$, and all paths are assembled into the final SVG file $S$. This output is directly compatible with standard vector graphics tools such as Adobe Illustrator.
Algorithm 1 Convert Bézier Splatting to Standard SVG.
  Input: Optimized control points $\mathcal{C}\in\mathbb{R}^{N\times(3k+1)\times 2}$, feature colors $\mathcal{F}\in\mathbb{R}^{N\times 3}$, canvas size $(W,H)$
  Output: SVG file $S$ containing cubic Bézier paths
  for each curve $i=1$ to $N$ do
     $P_{i}\leftarrow\frac{\mathcal{C}_{i}+1}{2}\cdot(W,H)$
     $(r,g,b)\leftarrow\text{sigmoid}(\mathcal{F}_{i})\times 255$
     Initialize path $d\leftarrow\texttt{[M $P_{i}^{0}$]}$
     for each segment $j=1$ to $k$ do
        $(p_{1},p_{2},p_{3})\leftarrow$($P_{i}^{3j+1},P_{i}^{3j+2},P_{i}^{3j+3})$
        $d\oplus\texttt{[C $p_{1}$, $p_{2}$, $p_{3}$]}$
     end for
     $d\oplus\texttt{Z}$
     $\mathcal{S}\leftarrow\mathcal{S}\cup\{\text{path}(d,\,\text{fill}=(r,g,b),\,\text{opacity}=1.0)\}$
  end for
  Return: SVG file assembled from all paths
Appendix C Bézier Splatting Supports Flexible Attributes of Curves
All 2D Gaussians in our method are generated via a differentiable sampling algorithm, which makes it straightforward to incorporate user-defined shape attributes, such as linear-gradient fills in color or opacity. To demonstrate the extensibility of our model, we evaluate the multi-opacity scheme for open curves: each Bézier segment within an open curve is assigned an independent opacity value. This design enables users to either maintain consistent sampling point appearance across segments or apply customized interpolation strategies. This strategy significantly improves vectorization quality for open curves, as shown in Table  5. Since closed curves are the primary representation format in SVG, we adopt a uniform setting for opacity and color in those regions to ensure a full compatibility with existing vector graphic tools. Nonetheless, users can still follow the above multi-opacity scheme to implement linear-gradient fills for both opacity and color, enhancing the flexibility and expressiveness of vector graphics.
Table 5: An ablation study on the number of opacity parameters per curve, evaluated on 512 open curves with DIV2K dataset [23].
@@ Appendix C Bézier Splatting Supports Flexible Attributes of Curves
3-opacity 0.65 23.79 0.50
@@
Appendix D More Comparisons
We present more qualitative comparisons on the DIV2K [23] dataset. As shown in Fig. 9, Fig. 10, Fig. 11, Fig. 12, and Fig. 13, our method shows consistent improvements over the baselines, including better preservation of fine details, higher rendering fidelity, fewer visual artifacts, and more accurate geometric structures. Furthermore, as the number of curves increases, our method demonstrates a significantly improved representational ability, enabling more precise reconstruction of complex shapes, and fine-grained structures.
We further evaluate our method on another natural image dataset, Kodak [13]. Due to the slow vectorization speed of LIVE [17], we compare with DiffVG on open curves only. As shown in Table 6, our method consistently outperforms DiffVG across quantitative metrics including PSNR, SSIM, and LPIPS.
@@ Appendix D More Comparisons
Table 6: A quantitative evaluation on the Kodak dataset [13] with 256 to 1024 curves. We report results on open curves.
Method 256 512 1024
$SSIM^{\uparrow}$ $PSNR^{\uparrow}$ $LPIPS^{\downarrow}$ $SSIM^{\uparrow}$ $PSNR^{\uparrow}$ $LPIPS^{\downarrow}$ $SSIM^{\uparrow}$ $PSNR^{\uparrow}$ $LPIPS^{\downarrow}$
Open DiffVG 0.601 23.43 0.535 0.645 24.70 0.495 0.699 26.04 0.439
Ours 0.679 26.18 0.457 0.743 27.90 0.383 0.797 29.24 0.310
Closed DiffVG 0.622 24.11 0.513 0.666 25.34 0.475 0.719 26.66 0.420
Ours 0.621 24.19 0.519 0.664 25.61 0.485 0.708 26.91 0.448
Appendix E More Image Vectorization Results
@@
Fig. 14, Fig. 15, and Fig. 16 show more results on natural images from DIV2K [23] and Kodak [13], as well as animation images from DanbooRegion [31], respectively. The results demonstrate that both the global structure and local texture of the images are well reconstructed, highlighting the effectiveness of our approach in capturing fine details and complex shapes.
@@
Figure 10: More qualitative comparisons of our method and the existing differentiable VG rasterization method on DIV2K dataset [23].
Figure 11: More qualitative comparisons of our method and the existing differentiable VG rasterization method on DIV2K dataset [23].
Figure 12: More qualitative comparisons of our method and the existing differentiable VG rasterization method on DIV2K dataset [23].
Figure 13: More qualitative comparisons of our method and the existing differentiable VG rasterization method on DIV2K dataset [23].
Figure 14: More image vectorization results by Bézier Splatting on DIV2K dataset [23].
Figure 15: More image vectorization results by Bézier Splatting on Kodak [13] dataset.
Figure 16: More image vectorization results by Bézier Splatting on DanbooRegion dataset [31].